OpenStudy (anonymous):

solve for X. log(base3)x^2=2log(base3)4-4log(base3)5

OpenStudy (anonymous):

like this? \[\log_3(x^2) = 2\log_3(4)-4\log_3(5)\]

OpenStudy (anonymous):

log_3(x^2)=2log_3(4)-4log_3(5)??????

OpenStudy (anonymous):

x^2 = 3^RHS x = sqrt(3^RHS) I think?

OpenStudy (anonymous):

x = 0.16

OpenStudy (anonymous):

since it is all base three, lets simplify each side down to one log. The left side is already done. On the right side we have: \[2\log_3 (4)-4\log_3(5) = \log_3(4^2)-\log_3(5^4) = \log_3\left(\frac{16}{625}\right)\] So now we have: \[\log_3(x^2) = \log_3\left(\frac{16}{625}\right) \iff x^2 = \frac{16}{625} x = \pm \frac{4}{25}\]

OpenStudy (anonymous):

whoops there should be an arrow in between the fraction 16/625 and x in the last line.